OpenStudy (anonymous):

2n^2 -5n -12 my helper is not around can someone helps me SOLVE THIS

OpenStudy (anonymous):

COULD THIS START WID 2N

OpenStudy (anonymous):

\[(2n + a)(n + b) = 2n^2 - 5n - 12\]\[\implies2b + a = -5\]and\[a*b = -12\]

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

UMMMM 3 AND 4

OpenStudy (anonymous):

Right, but one must be negative and one must be positive if they multiply to make -12.

OpenStudy (anonymous):

And since they add to make -5, probably a good bet that it's the bigger one that's negative.

OpenStudy (anonymous):

-4 , +3

OpenStudy (anonymous):

4X^2 -36X + 30 I KNOW ITS START WITH 2X AND ALSO IT HAS 15 N 6 BUT I DONT KNOW TO PUT THEM TOGTHER

OpenStudy (anonymous):

Are you sure you have that right? That trinomial doesn't factor well.

OpenStudy (anonymous):

4X^2 -26X + 30

OpenStudy (anonymous):

That's better. I think you'll find that the factors have the form: (4x + a)(x + b) Not: (2x + a)(2x + b)

OpenStudy (anonymous):

So that means that 4b + a = -26 and a * b = 30

OpenStudy (anonymous):

5 N 6